jfruge
New member
Well lamp column 2 is clearly not working for ya. Check the path of the YELLOW-RED wire from the laps back to the header 1J7-2. With the game off, use your meter in continuity mode to determine if there is an opening anywhere along the YELLOW-RED wire path to the 1J7-2 header. Might be a good idea to reflow the 1J6 and 1J7 connectors while you're at it.
For a quick test you could also swap the PIAs at U14 and U15 to see if the problem follows. If nothing changes then you'll need to work from U14-11 through U54-12->11, Q16 and Q24.
You're lucky the problem is so consistent and well defined.
Ok, I will try the trouble shooting steps you called out
Thanks,